OpenStudy (anonymous):

Car with mass 940kg is moving at speed 48km/h. How much force do you need to stop it in 15m distance?

OpenStudy (ghazi):

first of all find the retardation \[(48)^2=2*a*D\] and \[a=\frac{ (48)^2 }{ 2*15 }=76.8m/s^2\] now F=m*a F=940*76.8 N @duck123 you are done

OpenStudy (ash2326):

@duck123 Did you understand?

OpenStudy (anonymous):

I'm trying to figure out where 2*a*d part is coming from

OpenStudy (ghazi):

\[V^2=U^2-2a*D\] V is final velocity which is zero and U is = 48 i have taken -a because it is retardation hope that helps

OpenStudy (anonymous):

oh, now i get it. thanks!

OpenStudy (ghazi):

:)